OpenStudy (anonymous):

what number is 42% of 150

OpenStudy (anonymous):

42%=.42 150*.42=63

OpenStudy (e.cociuba):

42 percent of 150 is 0.42*150 or 63. And conversely, 150% of 42 is 63 as well.

OpenStudy (mathstudent55):

Translate English sentence into math sentence (also known as an equation): Unknown number = x 42 = 42 % = 1/100 = 0.01, so 42% = 0.42 "of" = "times" = * 150 = 150 So, what number is 42% of 150 becomes x = 0.42 * 150 Just multiply 0.42 and 150 together.